OpenStudy (anonymous):

Derek gave Kyle half of his marble collection. Kyle gave Janis half of the marbles he received from Derek. Janis kept 12 of the marbles and gave the remaining 16 to Jill. How many marbles did Derek give to Kyle?

OpenStudy (anonymous):

Janis kept 12 of the marbles and gave the remaining 16 to Jill. so janis got 12+16=18 kyle got=28*2=56

OpenStudy (anonymous):

:) nice